OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 120r3 square inches. Each tile has an area of square root of the quantity 16 r to the ninth power. The total number of tiles used can be represented by the expression below. one hundred twenty r to the third power, all over the square root of the quantity sixteen r to the ninth power Simplify the expression for the total number of tiles used. Show your work. @Nurali @zepdrix

zepdrix (zepdrix):

\[\Large A=120r^3 \qquad\text{Total Area}\]\[\Large a=16r^9 \qquad\text{Area of each tile}\] \[\Large n=\frac{A}{a} \qquad\text{Number of tiles}\] \[\Large n=\frac{120r^3}{16r^9}\] So let's divide the constants and the r's separately.\[\Large n=\frac{120}{16}\cdot\frac{r^3}{r^9}\]What does the division of the constants give us?

zepdrix (zepdrix):

120/16 = ?

OpenStudy (anonymous):

One moment please, I have to write all of this on a program, and it takes a bit of time.

zepdrix (zepdrix):

Oh I see :) lol

OpenStudy (anonymous):

Like this: http://www.sketchtoy.com/48555621 and 120/16 is 7.5

zepdrix (zepdrix):

\[\Large n=7.5\cdot\frac{r^3}{r^9}\]Ok good. To divide exponentials of the same base, we `subtract` the exponents. Example:\[\Large \frac{x^\color{green}{2}}{x^\color{royalblue}{5}}=x^{\color{green}{2}-\color{royalblue}{5}}=x^{-3}\]

OpenStudy (nurali):

http://www.sketchtoy.com/48555862

OpenStudy (anonymous):

@zepdrix x^-6

zepdrix (zepdrix):

r^-6 for our problem? Yay good job :)

OpenStudy (anonymous):

Is that all @zepdrix ?

zepdrix (zepdrix):

\[\Large n=7.5r^{-6}\]Yes. You could rewrite it as Nurali did to make it look a bit prettier,\[\Large n=\frac{15}{2r^6}\]But it's no big deal :)

OpenStudy (anonymous):

How did Nurali get that? o.o

zepdrix (zepdrix):

\[\Large 7.5\cdot\frac{2}{2}=\frac{15}{2}\]That was to deal with the decimal.

OpenStudy (anonymous):

Oh, okay.

zepdrix (zepdrix):

For the other part, you need to recall a rule of exponents. We can switch the sign on the exponent by tossing it into the denominator.\[\Large r^{-6}=\frac{1}{r^6}\]

OpenStudy (anonymous):

Thanks again! ^_^

zepdrix (zepdrix):

yay team \c:/